# Differentially private and decentralized randomized power method

### v2

@@ A.2.1 Note on related privacy proofs:
We can see lines 4-5 of Algorithm 1 as a sequential composition ($M$) of $L$ Gaussian Mechanisms. By Lemma 2.1, each mechanism $M_{i}$ satisfies ($\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP where $\Delta_{i}$ is the $\ell_{2}$-sensitivity of the function associated to mechanism $\Delta_{i}$ and $\sigma_{i}^{2}$ is the variance of the noise added with the Gaussian Mechanism. By Lemma 2.2, the composition of mechanisms $M=(M_{1},...,M_{i},...,M_{L})$ satisfies ($\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP. Let $\rho\triangleq\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$. By design of our algorithm, we have:
@@
By Proposition 1.3 of Bun & Steinke (2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$. Let $\epsilon^{\prime}\triangleq\rho+2\sqrt{\rho\log(1/\delta)}$.
Then:
@@
which is a reasonable assumption, as in practice $\epsilon=O(1)$ and $\delta\ll\frac{1}{d}$ where $d$ is the number of records to protect. In our case, $d=n^{2}$ because we run the privacy-preserving power method on $\bm{A}\in\mathbb{R}^{n\times n}$. ∎
@@
According to (Hardt & Price, 2014), if $\bm{G}_{l}\sim\mathcal{N}(0,\sigma_{l})^{d\times p}$, then with probability p=99/100 we have:
@@
We can therefore satisfy the noise conditions of Theorem 3.3 with probability $p$ if we choose $\eta=\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}$, giving us:
$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\cdot\bm{U}_{k}\|\leq\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}\,.$ (41)
Which leads us to the following Theorem by injecting $\sigma_{l}=\hat{\Delta}_{l}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$.
A.4 Derivation of Approximation 6.2
We provide here the derivation for Approximation 6.2:
Proof.
Let $\bm{M}\in\mathbb{R}^{n\times k}$ with i.i.d $\mathcal{N}(0,1)$-distributed entries. Let $\bm{M}=\bm{Q}\bm{R}$ be its QR factorization (by definition $\bm{i}sQ$ orthogonal and $\bm{R}$ upper triangular). By the Barlett decomposition theorem (Muirhead, 2009), we know that $\bm{Q}$ is a random matrix distributed uniformly in the Stiefel manifold $\mathbb{V}_{k,n}$. Then by Theorem 2.2.1 of Chikuse (2012), we know that a random matrix $\bm{Q}$ uniformly distributed on $\mathbb{V}_{k,n}$ can be expressed as $\bm{Q}=\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}}$ with $\bm{Z}$ another matrix with i.i.d $\mathcal{N}(0,1)$-distributed entries. We approximate $\bm{Z}^{\top}\bm{Z}$ as a diagonal matrix and remark that its diagonal elements are distributed as random chi-squared variables with $n$ degrees of liberty. For a matrix $\bm{A}$, we note as $\bm{A}^{\circ n}$ the elementwise Hadamard exponentiation. Then we have:
$\bm{Q}_{ij}^{2}$ $=(\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})_{ij}^{2}$ (42)
$=(\bm{Z}^{\circ 2}((\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})^{\circ 2})_{ij}$ (43)
$=(\bm{Z}^{\circ 2}(\bm{Z}^{\top}\bm{Z})^{-1})_{ij}.$ (44)
We remark that each element of $\bm{Z}^{\circ 2}$ is distributed as a chi-squared variable with one degree of freedom, therefore $n\cdot\bm{Q}_{ij}^{2}$ is distributed as an i.i.d $F(1,n)$ random variable by definition. For large $n$, we can approximate these $F$ variables as Gaussians with mean $\mu=\frac{n}{n-2}$ and variances $\sigma^{2}=\frac{2n^{2}(n-1)}{(n-2)^{2}(n-4)}$. Therefore $\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{\mu}{n},\frac{\sigma^{2}}{n^{2}})$ and:
$k\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k^{2}\sigma^{2}}{n^{2}})$ $\sum_{j}\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k\sigma^{2}}{n^{2}}).$ (45)
We can approximate the expectancy of the maximum (noted as m) of $d$ Gaussian variables distributed as $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ by $m=\sigma_{2}\sqrt{2\cdot\log(d)}+\mu_{2}$ by Lemma 2.3 from Massart (1896). We therefore get approximations of $\mathbb{E}[\max_{ij}k\bm{Q}_{ij}^{2}]$ and $\mathbb{E}[\max_{i}\|\bm{Q}_{i:}\|^{2}]$. We assume that the variances of $\Delta_{l}^{prior}$ and $\hat{\Delta}_{l}$ (as $n$ and $k$ are large) are small and therefore:
$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\mathbb{E}[(\Delta_{l}^{prior})^{2}]=\sqrt{\frac{k^{2}\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(kn)}+\frac{k\mu}{n}$ (46)
Figure 3: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=12000$ and $k$ ranging between 64 and 4000, with a step of 64. Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm. The theoretical ratios are estimated using Theorem 6.2.
A.5 Proof of Theorem 4.1
@@
First bound: We know by Section 2.4 from Woodruff et al. (2014) that any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$ has squared norm at most 1. Therefore $\max_{i}\|\bm{X}_{i:}\|_{2}\leq 1$ because $\bm{X}$ has orthonormal columns by construction.
Second bound: Without loss of generality and to simplify notation, we note $\bm{X}$ to denote a matrix $\bm{X}_{l}$ computed during the PPM. As proposed in Hardt & Price (2014), we can write any column $x_{c}$ of $\bm{X}$ ($\bm{X}_{:c}$) as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$ where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar. Therefore we have:
$<x_{c},x_{o}>$ $=<\sum_{i=1}^{n}\alpha_{c}^{i}u_{i},\sum_{j=1}^{n}\alpha_{o}^{i}u_{j}>$ (48)
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}<u_{i},u_{j}>$ (By bilinearity) (49)
$=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}.$ (Orth. columns of $\bm{U}$) (50)
$\bm{X}$ has orthonormal columns, so if $c=o$, then $<x_{c},x_{o}>=\sum_{i=1}^{n}{(\alpha_{c}^{i})}^{2}=1$. Otherwise, we have $<x_{c},x_{o}>=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}=0$. We can therefore define a matrix $\bm{B}$ with orthonormal columns such that $\bm{X}=\bm{U}\bm{B}$ and $\bm{B}_{jc}=\alpha_{c}^{j}$. We can therefore write any row $i$ of $\bm{X}$ as $\bm{X}_{i:}=\sum_{j=1}^{p}\bm{U}_{ij}\bm{B}_{j:}$. Therefore
$\|\bm{X}_{i:}\|_{2}$ $=\|\sum_{j=1}^{p}\bm{U}_{ij}\bm{B}_{j:}\|_{2}$ (51)
$\leq\sum_{j=1}^{p}\|\bm{U}_{ij}\bm{B}_{j:}\|_{2}$ (52)
$\leq\sum_{j=1}^{p}|\bm{U}_{ij}|\|\bm{B}_{j:}\|_{2}$ (53)
$\leq\sum_{j=1}^{p}|\bm{U}_{ij}|.$ (Orth. columns of $\bm{B}$) (54)
We then have $\max_{i}\|\bm{X}_{i:}\|_{2}\leq\max_{i}(\sum_{j=1}^{p}|\bm{U}_{ij}|)=\|\bm{U}\|_{\infty}=\mu_{1}(\bm{A})$.
We proposed two bounds for $\|\bm{X}_{i:}\|_{2}$, which are useful in two different cases. We can however note that:
$\|\bm{X}_{i:}\|_{2}\leq 1\text{ and }\|\bm{X}_{i:}\|_{2}\leq\mu_{1}(\bm{A})$ (55)
$\Rightarrow\|\bm{X}_{i:}\|_{2}\leq\min(1,\mu_{1}(\bm{A})),$ (56)
which leads us to Theorem 4.1 giving us a runtime-independent bound.
A.6 Proof of Theorem 5.2
@@
A.7 Application to recommender systems:
@@
To simplify our analysis, we consider $\bm{I}$ public and do not use item-wise normalization in the computation of ideal low pass filter, leaving it for future work. We therefore define $\tilde{\bm{R}}=\bm{U}^{-\frac{1}{2}}\bm{R}$ and $\tilde{\bm{P}}=(\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})$.
Lemma A.1.
@@
We note $\tilde{\bm{P}}_{ij}$ as the element of $\tilde{\bm{P}}$ at line $i$ and column $j$ and $d_{user}(u)=\sum_{i=0}^{n-1}r_{ui}$. We can write $\tilde{\bm{P}}_{ij}$ as such:
$\begin{split}\tilde{\bm{P}}_{ij}&=((\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R}))_{ij}\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R})^{T})_{i,*}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=((\bm{U}^{-\frac{1}{2}}\bm{R})_{*,i})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=\sum_{u=0}^{s-1}\frac{1}{\sqrt{d_{user}(u)}}r_{ui}\frac{1}{\sqrt{d_{user}(u)}}r_{uj}\\ &=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)}\cdot r_{ui}\cdot r_{uj}.\\ \end{split}$ (59)
By noticing that $(\bm{R}^{\top}\bm{R})_{ij}=\sum_{u}r_{ui}\cdot r_{uj}$, then we can deduce that $\tilde{\bm{P}}=\sum_{u}\frac{1}{d_{user}(u)}\bm{R}_{u}^{\top}\bm{R}_{u}$. Therefore $\tilde{\bm{P}}$ is partitioned among $s$ users as described in Section 5. Sensitivity: We protect the user at the item-level and use the deletion model of differential privacy to compute the sensitivity of the PPM used with the item-item normalized adjacency matrix ($\bm{A}=\tilde{\bm{P}}$). Therefore we have:
@@
where $r_{ui}^{\prime}=r_{ui}$ except for one user-item interaction, i.e. $r_{vk}=1$ but $r_{vk}^{\prime}=0$. Let $\bm{C}=\bm{A}-\bm{A}^{\prime}$. Let $\mathcal{N}(v)$ be the set of items which user $v$ liked before deletion. We have:
@@
$+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}.$ (66)
@@
$+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (80)
@@
A.7.1 Approximation errors comparisons:
As explained in Section 6.1.2, GF-CF and BSPM (with its parameter $T_{b}=1$) use $\bm{U}_{p}$ to compute the ideal low pass filter and filter the interaction matrix $\bm{R}$, yielding $\bm{R}_{p}$. Indeed, we have:

### v3

@@ A.2.2 Zero-Concentrated Differential Privacy (zCDP):
We can see lines 4-5 of Algorithm 1 as a sequential composition ($M$) of $L$ Gaussian Mechanisms. By Lemma A.1, each mechanism $M_{i}$ satisfies ($\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP where $\Delta_{i}$ is the $\ell_{2}$-sensitivity of the function associated to mechanism $\Delta_{i}$ and $\sigma_{i}^{2}$ is the variance of the noise added with the Gaussian Mechanism. By Lemma A.2, the composition of mechanisms $M=(M_{1},...,M_{i},...,M_{L})$ satisfies ($\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP. Let $\rho\triangleq\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$. By design of our algorithm, we have:
@@
By Proposition 1.3 of Bun and Steinke (2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$. Let $\epsilon^{\prime}\triangleq\rho+2\sqrt{\rho\log(1/\delta)}$. Then:
@@
which is a reasonable assumption, as in practice $\epsilon=O(1)$ and $\delta\ll\frac{1}{d}$, where $d$ is the number of records to protect. In our case, $d=n^{2}$ because we run the privacy-preserving power method on $\bm{A}\in\mathbb{R}^{n\times n}$. ∎
@@
The following theorem from Balcan et al. (2016) will be useful in our proof:
Theorem A.3 (Bound for the noisy power method (NPM) (Balcan et al., 2016)).
Let $k\leq q\leq p$ be positive integers. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of a positive semi-definite matrix $\bm{A}$ and let $\lambda_{1}\geq\cdots\geq\lambda_{d}\geq 0$ denote its eigenvalues and fix $\eta=O\left(\frac{\lambda_{q}}{\lambda_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\lambda_{k}}{\lambda_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right)$. If at every iteration $l$ of the NPM $\bm{G}_{\ell}$ satisfies Assumption 2.1, then after
$L=\Theta\left(\frac{\lambda_{k}}{\lambda_{k}-\lambda_{q+1}}\log\left(\frac{\tau d}{\eta}\right)\right).$
$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\bm{U}_{k}\|_{2}\leq\eta\quad\text{and}\quad\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{A}\|_{2}^{2}\leq\lambda_{k+1}^{2}+\eta^{2}\lambda_{k}^{2}.$ (32)
We now provide the proof of Theorem 3.3:
@@
According to Hardt and Price (2014), if $\bm{G}_{l}\sim\mathcal{N}(0,\sigma_{l})^{d\times p}$, then with probability 99/100 we have:
@@
We can therefore satisfy the noise conditions of Theorem A.3 with probability 99/100 if we choose $\eta=\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\lambda_{k}-\lambda_{q+1}}$, giving us:
$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\cdot\bm{U}_{k}\|\leq\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\lambda_{k}-\lambda_{q+1}}\,,$ (34)
which leads us to the statement of the theorem by injecting $\sigma_{l}=\hat{\Delta}_{l}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$. ∎
A.4 Proof of Theorem 4.1
@@
We showed in Equation that $\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}$. This quantity depends on values computed during the execution of the algorithm. We now show that we can bound $\max_{i}\|\bm{X}_{i:}^{l}\|_{F}$ with a runtime-independent bound.
Without loss of generality and to simplify notation, we use $\bm{X}$ to denote any matrix $\bm{X}_{l}$ computed during the execution of the Private Power Method. Let $\bm{X}_{:c}=x_{c}$ denote a column of $\bm{X}$.
As $\bm{A}$ is an Hermitian matrix, by the spectral theorem, we have $\bm{A}=\bm{U}\bm{D}\bm{U}^{\top}$, where $\bm{U}$ is unitary (with orthonormal columns) and $\bm{D}$ is diagonal.
As the columns of $\bm{U}$ form a complete basis for $\mathbb{R}^{n}$, we can write any column $x_{c}$ of $\bm{X}$ as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$, where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar.
We can then write:
$\langle x_{c},x_{e}\rangle$ $=\langle\sum_{i=1}^{n}\alpha_{c}^{i}u_{i},\sum_{j=1}^{n}\alpha_{e}^{j}u_{j}\rangle$ (35)
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{c}^{i}\alpha_{e}^{j}\langle u_{i},u_{j}\rangle$ (36)
$=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{e}^{i}.$ (Orth. columns of $\bm{U}$) (37)
$\bm{X}$ is the “$Q$“ matrix constructed from a Gram-Schmidt QR decomposition, it has therefore orthonormal columns by definition. Therefore, if $c=e$, then $\langle x_{c},x_{e}\rangle=\sum_{i=1}^{n}{(\alpha_{c}^{i})}^{2}=1$. Otherwise, we have $\langle x_{c},x_{e}\rangle=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{e}^{i}=0$.
The key is then to notice that we can define a matrix $\bm{B}\in\mathbb{R}^{n\times p}$ with orthonormal columns such that $\bm{X}=\bm{U}\bm{B}$ and $\bm{B}_{jc}=\alpha_{c}^{j}$.
We recall that any matrices $\bm{H}$ and $\bm{J}$, we have $\|\bm{H}\bm{J}\|_{F}\leq\|\bm{H}\|_{2}\|\bm{J}\|_{F}$.
Then, we can bound the norm of any row of $\bm{X}_{i:}$ as:
$\|\bm{X}_{i:}\|_{F}$ $=\|\bm{X}_{i:}^{\top}\|_{F}\,,$ (38)
$=\|\bm{B}^{\top}\bm{U}_{i:}^{\top}\|_{F}\,,$ (39)
$\leq\|\bm{B}^{\top}\|_{2}\|\bm{U}_{i:}^{\top}\|_{F}.$ (40)
$\bm{B}$ has orthonormal columns by construction, therefore
$\|\bm{B}\|_{2}$ $=\sqrt{\|\bm{B}^{\top}\bm{B}\|_{2}}$ (41)
$=\sqrt{\|\bm{I}\|_{2}}$ (42)
$=1.$ (43)
We then have:
$\|\bm{X}_{i:}\|_{F}$ $\leq\|\bm{B}^{\top}\|_{2}\|\bm{U}_{i:}^{\top}\|_{F}$ (44)
$\leq\|\bm{U}_{i:}^{\top}\|_{F}$ (45)
$\leq\|\bm{U}_{i:}\|_{F}.$ (46)
Additionally as $\bm{X}_{i:}\in\mathbb{R}^{1\times n}$, $\max_{i}\|\bm{X}_{i:}\|_{2}=\max_{i}\|\bm{X}_{i:}\|_{F}\leq\max_{i}(\|\bm{U}_{i:}\|_{F})=\mu_{1}(\bm{A})$.
Note: By Section 2.4 from Woodruff (2014), for any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$, $\|v\|_{2}\leq 1$.
As $\bm{U}$ has orthonormal columns by construction, $\max_{i}\|\bm{U}_{i:}\|_{2}\leq 1$.
We can therefore bound $\max_{i}\|\bm{X}_{i:}\|_{2}$ as:
$\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}\|_{2}\leq\mu_{1}(\bm{A})\leq 1.$ (47)
Injecting this bound in Theorem 3.3 leads us to Theorem 4.1, giving us a runtime-independent bound. ∎
A.5 Proof of Theorem 5.2
@@
Appendix B Application to recommender systems:
@@
To simplify our analysis, we consider $\bm{I}$ public and do not use item-wise normalization in the computation of the ideal low pass filter, leaving it for future work. We therefore define $\tilde{\bm{R}}=\bm{U}^{-\frac{1}{2}}\bm{R}$ and $\tilde{\bm{P}}=(\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})$.
Lemma B.1.
@@
Let $\tilde{\bm{P}}_{ij}$ denote the element of $\tilde{\bm{P}}$ at row $i$ and column $j$ and let $d_{user}(u)=\sum_{i=0}^{n-1}r_{ui}$. We can write $\tilde{\bm{P}}_{ij}$ as:
$\begin{split}\tilde{\bm{P}}_{ij}&=((\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R}))_{ij}\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R})^{\top})_{i,*}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=((\bm{U}^{-\frac{1}{2}}\bm{R})_{*,i})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=\sum_{u=0}^{s-1}\frac{1}{\sqrt{d_{user}(u)}}r_{ui}\frac{1}{\sqrt{d_{user}(u)}}r_{uj}\\ &=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)}\cdot r_{ui}\cdot r_{uj}.\\ \end{split}$ (50)
By noticing that $(\bm{R}^{\top}\bm{R})_{ij}=\sum_{u}r_{ui}\cdot r_{uj}$, we can deduce that $\tilde{\bm{P}}=\sum_{u}\frac{1}{d_{user}(u)}\bm{R}_{u}^{\top}\bm{R}_{u}$. Therefore $\tilde{\bm{P}}$ is partitioned among $s$ users as described in Section 5.
Sensitivity: We protect the user at the item-level and use the deletion model of differential privacy to compute the sensitivity of the PPM used with the item-item normalized adjacency matrix ($\bm{A}=\tilde{\bm{P}}$). Therefore we have:
@@
where $r_{ui}^{\prime}=r_{ui}$ except for one user-item interaction, i.e., $r_{vk}=1$ but $r_{vk}^{\prime}=0$. Let $\bm{C}=\bm{A}-\bm{A}^{\prime}$. Let $\mathcal{N}(v)$ be the set of items which user $v$ liked before deletion. We have:
@@
$\quad\quad+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}.$ (57)
@@
$\quad\quad+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (68)
@@
B.1 Approximation errors comparisons:
As explained in Section 6.1, GF-CF and BSPM (with its parameter $T_{b}=1$) use $\bm{U}_{p}$ to compute the ideal low pass filter and filter the interaction matrix $\bm{R}$, yielding $\bm{R}_{p}$. Indeed, we have: